# Université Paris Cité, F-75005 Paris, France

### v1

@@ 1 Introduction
where $Q_{i,n}$ is the value of the charge $Q_{i}$ in state $n$. For a review of the role of GGEs in the contexts of statistical mechanics and thermalisation see DAlessio:2015qtq .
@@
Often it is not enough for our theory to have an infinite set of conserved charges, we also want the charges to commute. It has been known for some time that 2d CFTs contain infinite sets of mutually commuting conserved charges Sasaki:1987mm . The most well known set of charges are related to the classical KdV hierarchy, as detailed in Bazhanov:1994ft , and hence we will refer to them as the KdV charges. They are constructed from the Virasoro modes and we list the first three here
@@
These GGEs have been studied extensively in the literature. Their large central charge limit ($c\rightarrow\infty$) was studied in a series of papers by Dymarsky et al Dymarsky:2018lhf ; Dymarsky:2018iwx ; Dymarsky:2022dhi and also by Maloney et al in Maloney:2018yrz . There, expressions for these GGEs in the limit $c\rightarrow\infty$ and leading $1/c$ corrections were derived. These GGEs are then holographically dual to a class of black holes in AdS3 referred to in Dymarsky:2020tjh as KdV charged black holes and their connection to the eigenstate thermalisation hypothesis was also explored in Dymarsky:2019etq .
In this paper we will be investigating the modular properties of these GGEs. The deep relationship between 2d CFTs and modular forms has been known for a long time and has been used extensively to study 2d CFTs. It is know that the characters of a rational 2d CFT form a vector valued modular form. This was first suggested by Cardy in CARDY1986186 and rigorously proven by Zhu in Zhu1996 . It is also known that if we expand the GGE (6) as a power series in chemical potentials $\alpha_{2n-1}$, then each term, which is a correlation functions of the charges, is a modular form or quasi-modular form. This was first argued by Dijkgraaf in Dijkgraaf:1996iy and then in Maloney:2018hdg Maloney et al found expressions for the correlators in terms of modular differential operators acting on the characters which makes the modular properties manifest.
A natural question to ask is whether the full GGE has any interesting modular properties. This has been studied in detail for the GGEs in the free fermion model ($c=\frac{1}{2}$ Ising minimal model) in the series of papers Downing:2021mfw ; Downing:2023lnp ; Downing:2023lop . In general, closed form expressions for the GGEs are not known. However for the free fermion model the simplicity of the theory means that exact expressions can be found and used as a starting point to study the modular properties. This meant an explicit expression for the modular transformation could be found.
@@
However an exact modular transform can be found. This was done by using the thermodynamic Bethe ansatz (TBA). The TBA for the original GGE is known from Bazhanov:1996aq . We then take a mirror transform of the TBA in order to find the spectrum of the GGE in the new channel. When this is done the spectrum from the asymptotic results can be reproduced, but we also find additional energies. These energies behave as $C\alpha^{-\nu}$, where $\alpha<0$ is the chemical potential, Re$(C)>0$ and $\nu>0$, and hence when they are exponentiated they give rise to terms that have a vanishing asymptotic expansion. This is why they were missed in the original asymptotic analysis but including them in the transformation gives an exact expression for the modular transformed GGE.
@@
We then turn our attention to the TBA in section 5. We start by using the TBA to reproduce the asymptotic results. We then show that there are other solutions to the TBA equations which when exponentiated have a vanishing asymptotic expansion, just as in the case of the free fermion model. We conjecture that including these additional energies in the transformed GGE will give the exact modular transformation for the GGE. During this process we derive new integral equations that encode the spectrum of the KdV charges as well as the charges coming from the other quasi-primary fields in the theory.
@@ 2 Transformed GGEs and defects
The GGE can be thought of as the insertion of a defect as was done in Downing:2023lnp . We consider our theory to be living on a cylinder of circumference $R$ and length $L$ as shown in diagram (I) of figure 1. We identify the ends of the cylinder so it becomes a torus with modular parameter $\hat{\tau}=iL/R$. The insertion of the KdV charge $I_{2n-1}$ is given by a horizontal defect wrapping the cylinder. The defect operator is $\hat{D}=e^{-L\alpha I_{2n-1}(R)}$ and the GGE is given by
@@
When $\alpha=0$ the defect isn’t present and the GGEs (7) are the characters of the 2d CFT. It is known that the characters form vector valued modular forms Zhu1996
@@
We will start with an asymptotic analysis, as $\alpha\rightarrow 0$, of the modular transformation of (7) in sections 3 and 4. There, as was also done in the asymptotic analysis in Downing:2021mfw , we will assume that the defect Hilbert space is just the irreducible representations of the Virasoro algebra. In Downing:2021mfw , where the free fermion model was studied, it was found that the defect Hamiltonian $H_{D}(L)$ had an asymptotic expansion as a sum over the other KdV charges
@@
While we have an asymptotic expression for the Hamiltonian $H_{D}(L)$, based on the results in Downing:2021mfw ; Downing:2023lnp ; Downing:2023lop we believe this is not the full picture. There, additional terms had to be added to the transformed spectrum that behaved as $\alpha^{-\nu}$ for $\nu>0$. These terms were missed in the asymptotic analysis since the exponential $e^{-\alpha^{-\nu}}$ has a vanishing asymptotic expansion as $\alpha\rightarrow 0$ from above. These additional terms are found in section 5.6 where the power $\nu$ is derived.
These additional terms that needed to be added to the transformed spectrum were determined in Downing:2021mfw ; Downing:2023lnp by using the thermodynamic Bethe ansatz (TBA). In section 5 we again use the TBA to find additional terms that we believe should be added to the transformed spectrum in order to give the full modular transformed GGE (12). These additional terms in the spectrum come from additional terms that have been added to the Hilbert space $\mathcal{H}_{D}$, hence this Hilbert space is no longer an irreducible representation of the Virasoro algebra.
@@ 3 GGEs in a Generic 2d CFT
We start by considering GGEs with a KdV charge inserted for a generic 2d CFT. For simplicity we will just consider inserting a single charge but this will already lead to interesting results. As was done in Downing:2021mfw , we will start by expanding the GGE as an asymptotic series in the chemical potential associated to the inserted charge. We can then modular transform each term using the results from Maloney:2018hdg . When this was done in Downing:2021mfw for the free fermion model ($c=\frac{1}{2}$ Ising minimal model) we found that the transformed expressions could be written as correlators of the other KdV charges. In the case of a generic CFT we will find that the transformed expressions are instead given by correlators of all the charges from quasi-primary fields, not just the KdV charges.
@@
The modular properties of the thermal correlators $\langle I_{3}^{n}\rangle_{i}$ where studied by A. Maloney et al in Maloney:2018hdg . There they showed the correlators can be written as modular linear differential operators acting on the characters of the CFT. In particular up to order $\alpha^{2}$ we have the following expressions for the correlators
@@
We now want to take the modular transform of each term in the asymptotic expansion of the GGE. We will just take the $S:\hat{\tau}\mapsto\tau=-1/\hat{\tau}$ transform. The characters (28) of a 2d CFT form a weight 0 vector valued modular form Zhu1996 , so under the $S$ modular transform we have
@@
where we have removed the $R$ and $L$ dependence. As a function of $\hat{\tau}$, $\langle I_{2m-1}^{n}\rangle_{i}(\hat{\tau})$ is a vector valued quasi-modular form of weight $2mn$ and depth $n-1$. This was shown in Dijkgraaf:1996iy by considering contact terms between the currents that give rise to the charges. Hence we can write it in the form
@@
where $F_{2mn-2p}(\hat{\tau})$ is a weight $2mn-2p$ vector valued modular form, which can be written as a modular differential operator acting on the characters of the theory Maloney:2018hdg . We can then take the modular transform of each term in (44) to obtain
@@
To end this section we note that there are two interesting cases in which we can do away with the additional charge $J_{5}$ appearing in (37). The first is when the charges $I_{5}$ and $J_{5}$ correspond to states which only differ by a null state (and are hence proportional to one another). This happens when $c=\frac{1}{2}$, which is the Ising Model central charge. This fact was used in the series of papers Downing:2021mfw ; Downing:2023lnp ; Downing:2023lop which studied the modular properties of GGEs in the Ising model. The second case is when the central charge is $c=-2$. The integrability of the KdV equations at $c=-2$ was studied in DiFrancesco:1991xm , although it is not clear at the moment how one would study this is in the context of a GGE. The theory at this central charge is logarithmic, and so the GGE would involve taking traces over logarithmic modules. A review of logarithmic CFTs can be found in Creutzig:2013hma .
@@ 4 Asymptotic Analysis of the GGE in the Lee-Yang Model
We will now repeat the analysis from the previous section for the Lee-Yang theory. We have chosen this theory since it is arguably the simplest interacting 2d CFT with only two Virasoro representations, one with $h=0$ and the other with $h=-1/5$. The theory therefore has two characters and they satisfy a second order modular differential equation as detailed in Gaberdiel:2008pr . Using this second order differential equation allows us to simplify the expression for the correlators found in Maloney:2018hdg . We can then use these simplified expressions to compute more of these correlators than was done in Maloney:2018hdg . In particular we can compute to high enough order to check whether the fact that the additional charges (which come from the other quasi-primary fields) don’t commute with the KdV charges stops us being able to re-exponentiate the transformed expression. In the GGE studied below with $I_{5}(R)$ inserted the non-commutativity is first present when we transform the $\langle{I_{5}^{6}}\rangle$ term. We confirm that we can indeed still exponentiate the transformed expression to formally give an expression for the modular transforms of the original GGE as a new GGE with an infinite set of charges inserted.
In the Lee-Yang theory, the quasi-primary field that gives $I_{3}(R)$ is now a null state and hence the correlators containing $I_{3}$ vanish, as was proved in Maloney:2018hdg . The next simplest case for a GGE here is the ensemble with $I_{5}(R)$ inserted
@@
It is worth noting here that we did not necessarily need to use the MLDO expressions for the thermal correlators to calculate these transformations. We could have used the method developed in Gaberdiel:2012yb to calculate the transformed expressions of thermal correlation functions. This method was used in, for example, Iles:2014gra to calculate the transformations of $W_{3}$ characters in terms of zero-modes of known currents in the theory. The advantage of using the MLDO expressions comes from the fact that the map going from the currents in a 2d CFT to the thermal correlation functions of their zero-modes has a non-trivial kernel [^1] We would like to thank G. M. T. Watts for this observation.. That is, if we used the method previously mentioned, then we would not know a priori whether certain parts of that expression vanished.
@@
Given that the charges $I_{2n-1}(L)$ do not commute with the charges $J_{2n-1}(L)$, one question that may be asked is “Is this re-exponentiation a reasonable thing to do?”. It seems that the answer is yes, and the fact that these charges do not commute does not affect our ability to formally re-write the transformed GGE as another GGE. Let us take some time to elaborate on this point. When we expand the right hand side of (57), the ordering of the charges in the correlators matters since they do not commute. This will lead to the presence of correlators that contain the same charges in different orders and we need to ensure that all the necessary correlators are present when we take the modular transformations of the $\langle{I_{5}^{n}}\rangle$ in the original GGE.
@@
It is worth mentioning briefly that $\langle{I_{5}J_{9}I_{5}J_{9}}\rangle$ and $\langle{I_{5}^{2}J_{9}^{2}}\rangle$ cannot independently be written as modular linear differential operators (MLDOs) acting on the characters of the theory, but this particular linear combination presented above does have a representation as an MLDO acting on the characters. We suspect that if one carefully studies the contact terms between the relevant currents associated to these charges, as was done in Dijkgraaf:1996iy for a different model, then it may become clear that indeed these expectation values separately cannot be written as MLDOs, however we have not performed this analysis.
@@
While we have found an asymptotic expression for the transformed GGE, or rather an expression with the leading charges in the transformed GGE, (57), we don’t believe that these match as functions. Firstly the right hand side of (57) contains an infinite sum in the charges. It is not clear if this sum is convergent, indeed in the case of free fermions the equivalent sum over charges was not convergent and had to be regularised Downing:2021mfw . In the case of free fermions this regularisation introduced functions with a branch cut. Hence while the original GGE was real, the transformed expression was complex. This problem was resolved by introducing additional terms in the transformed expression that came from the thermodynamic Bethe ansatz (TBA). These additional terms made the transformed expression real. It was then proved in Downing:2023lop that these additional terms gave expressions that matched exactly, not just asymptotically. We will now use the TBA for the Lee-Yang model to first reproduce our asymptotic results, and then find additional terms that we believe should be included in the transformed expression for the GGE.

### v2

# Modular Properties of Generalised Gibbs Ensembles

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
where $Q_{i,n}$ is the value of the charge $Q_{i}$ in state $n$. For a review of the role of GGEs in the contexts of statistical mechanics and thermalisation see [1].
@@
Often it is not enough for our theory to have an infinite set of conserved charges, we also want the charges to commute. It has been known for some time that 2d CFTs contain infinite sets of mutually commuting conserved charges [2]. The most well known set of charges are related to the classical KdV hierarchy, as detailed in [3], and hence we will refer to them as the KdV charges. They are constructed from the Virasoro modes and we list the first three here
@@
These GGEs have been studied extensively in the literature. Their large central charge limit ($c\rightarrow\infty$) was studied in a series of papers by Dymarsky et al [4, 5, 6] and also by Maloney et al in [7] and Brehm and Das in [8]. There, expressions for these GGEs in the limit $c\rightarrow\infty$ and leading $1/c$ corrections were derived. These GGEs are then holographically dual to a class of black holes in AdS3 referred to in [9] as KdV charged black holes and their connection to the eigenstate thermalisation hypothesis was also explored in [10].
In this paper we will be investigating the modular properties of these GGEs. The deep relationship between 2d CFTs and modular forms has been known for a long time and has been used extensively to study 2d CFTs. It is know that the characters of a rational 2d CFT form a vector valued modular form. This was first suggested by Cardy in [11] and rigorously proven by Zhu in [12]. It is also known that if we expand the GGE (6) as a power series in chemical potentials $\alpha_{2n-1}$, then each term, which is a correlation functions of the charges, is a modular form or quasi-modular form. This was first argued by Dijkgraaf in [13] and then in [14] Maloney et al found expressions for the correlators in terms of modular differential operators acting on the characters which makes the modular properties manifest.
A natural question to ask is whether the full GGE has any interesting modular properties. This has been studied in detail for the GGEs in the free fermion model ($c=\frac{1}{2}$ Ising minimal model) in the series of papers [15, 16, 17]. In general, closed form expressions for the GGEs are not known. However for the free fermion model the simplicity of the theory means that exact expressions can be found and used as a starting point to study the modular properties. This meant an explicit expression for the modular transformation could be found.
@@
However an exact modular transform can be found. This was done by using the thermodynamic Bethe ansatz (TBA). The TBA for the original GGE is known from [18]. We then take a mirror transform of the TBA in order to find the spectrum of the GGE in the new channel. When this is done the spectrum from the asymptotic results can be reproduced, but we also find additional energies. These energies behave as $C\alpha^{-\nu}$, where $\alpha<0$ is the chemical potential, Re$(C)>0$ and $\nu>0$, and hence when they are exponentiated they give rise to terms that have a vanishing asymptotic expansion. This is why they were missed in the original asymptotic analysis but including them in the transformation gives an exact expression for the modular transformed GGE.
@@
We then turn our attention to the TBA in section 5. We start by using the TBA to reproduce the asymptotic results. We then show that there are other solutions to the TBA equations which when exponentiated have a vanishing asymptotic expansion, just as in the case of the free fermion model. We conjecture that including these additional energies in the transformed GGE will give the exact modular transformation for the GGE. During this process we derive new integral equations that encode the spectrum of the KdV charges as well as the charges coming from the other quasi-primary fields in the theory. An earlier derivation of the TBA for GGEs can be found in [19], and the observation that all conserved quantities should be understood in terms of driving terms in the TBA can be found [20] in the context of Quantum Quenches in Integrable Field theories.
@@ 2 Transformed GGEs and defects
The GGE can be thought of as the insertion of a defect as was done in [16]. We consider our theory to be living on a cylinder of circumference $R$ and length $L$ as shown in diagram (I) of figure 1. We identify the ends of the cylinder so it becomes a torus with modular parameter $\hat{\tau}=iL/R$. The insertion of the KdV charge $I_{2n-1}$ is given by a horizontal defect wrapping the cylinder. The defect operator is $\hat{D}=e^{-L\alpha I_{2n-1}(R)}$ and the GGE is given by
@@
When $\alpha=0$ the defect isn’t present and the GGEs (7) are the characters of the 2d CFT. It is known that the characters form vector valued modular forms [12]
@@
We will start with an asymptotic analysis, as $\alpha\rightarrow 0$, of the modular transformation of (7) in sections 3 and 4. There, as was also done in the asymptotic analysis in [15], we will assume that the defect Hilbert space is just the irreducible representations of the Virasoro algebra. In [15], where the free fermion model was studied, it was found that the defect Hamiltonian $H_{D}(L)$ had an asymptotic expansion as a sum over the other KdV charges
@@
While we have an asymptotic expression for the Hamiltonian $H_{D}(L)$, based on the results in [15, 16, 17] we believe this is not the full picture. There, additional terms had to be added to the transformed spectrum that behaved as $\alpha^{-\nu}$ for $\nu>0$. These terms were missed in the asymptotic analysis since the exponential $e^{-\alpha^{-\nu}}$ has a vanishing asymptotic expansion as $\alpha\rightarrow 0$ from above. These additional terms are found in section 5.6 where the power $\nu$ is derived.
These additional terms that needed to be added to the transformed spectrum were determined in [15, 16] by using the thermodynamic Bethe ansatz (TBA). In section 5 we again use the TBA to find additional terms that we believe should be added to the transformed spectrum in order to give the full modular transformed GGE (12). These additional terms in the spectrum come from additional terms that have been added to the Hilbert space $\mathcal{H}_{D}$, hence this Hilbert space is no longer an irreducible representation of the Virasoro algebra.
@@ 3 GGEs in a Generic 2d CFT
We start by considering GGEs with a KdV charge inserted for a generic 2d CFT. For simplicity we will just consider inserting a single charge but this will already lead to interesting results. As was done in [15], we will start by expanding the GGE as an asymptotic series in the chemical potential associated to the inserted charge. We can then modular transform each term using the results from [14]. When this was done in [15] for the free fermion model ($c=\frac{1}{2}$ Ising minimal model) we found that the transformed expressions could be written as correlators of the other KdV charges. In the case of a generic CFT we will find that the transformed expressions are instead given by correlators of all the charges from quasi-primary fields, not just the KdV charges.
@@
The modular properties of the thermal correlators $\langle I_{3}^{n}\rangle_{i}$ where studied by A. Maloney et al in [14]. There they showed the correlators can be written as modular linear differential operators acting on the characters of the CFT. In particular up to order $\alpha^{2}$ we have the following expressions for the correlators
@@
We now want to take the modular transform of each term in the asymptotic expansion of the GGE. We will just take the $S:\hat{\tau}\mapsto\tau=-1/\hat{\tau}$ transform. The characters (28) of a 2d CFT form a weight 0 vector valued modular form [12], so under the $S$ modular transform we have
@@
where we have removed the $R$ and $L$ dependence. As a function of $\hat{\tau}$, $\langle I_{2m-1}^{n}\rangle_{i}(\hat{\tau})$ is a vector valued quasi-modular form of weight $2mn$ and depth $n-1$. This was shown in [13] by considering contact terms between the currents that give rise to the charges. Hence we can write it in the form
@@
where $F_{2mn-2p}(\hat{\tau})$ is a weight $2mn-2p$ vector valued modular form, which can be written as a modular differential operator acting on the characters of the theory [14]. We can then take the modular transform of each term in (44) to obtain
@@
To end this section we note that there are two interesting cases in which we can do away with the additional charge $J_{5}$ appearing in (37). The first is when the charges $I_{5}$ and $J_{5}$ correspond to states which only differ by a null state (and are hence proportional to one another). This happens when $c=\frac{1}{2}$, which is the Ising Model central charge. This fact was used in the series of papers [15, 16, 17] which studied the modular properties of GGEs in the Ising model. The second case is when the central charge is $c=-2$. The integrability of the KdV equations at $c=-2$ was studied in [21], although it is not clear at the moment how one would study this is in the context of a GGE. The theory at this central charge is logarithmic, and so the GGE would involve taking traces over logarithmic modules. A review of logarithmic CFTs can be found in [22].
@@ 4 Asymptotic Analysis of the GGE in the Lee-Yang Model
We will now repeat the analysis from the previous section for the Lee-Yang theory. We have chosen this theory since it is arguably the simplest interacting 2d CFT with only two Virasoro representations, one with $h=0$ and the other with $h=-1/5$. The theory therefore has two characters and they satisfy a second order modular differential equation as detailed in [23]. Using this second order differential equation allows us to simplify the expression for the correlators found in [14]. We can then use these simplified expressions to compute more of these correlators than was done in [14]. In particular we can compute to high enough order to check whether the fact that the additional charges (which come from the other quasi-primary fields) don’t commute with the KdV charges stops us being able to re-exponentiate the transformed expression. In the GGE studied below with $I_{5}(R)$ inserted the non-commutativity is first present when we transform the $\langle{I_{5}^{6}}\rangle$ term. We confirm that we can indeed still exponentiate the transformed expression to formally give an expression for the modular transforms of the original GGE as a new GGE with an infinite set of charges inserted.
In the Lee-Yang theory, the quasi-primary field that gives $I_{3}(R)$ is now a null state and hence the correlators containing $I_{3}$ vanish, as was proved in [14]. The next simplest case for a GGE here is the ensemble with $I_{5}(R)$ inserted
@@
It is worth noting here that we did not necessarily need to use the MLDO expressions for the thermal correlators to calculate these transformations. We could have used the method developed in [24] to calculate the transformed expressions of thermal correlation functions. This method was used in, for example, [25] to calculate the transformations of $W_{3}$ characters in terms of zero-modes of known currents in the theory. The advantage of using the MLDO expressions comes from the fact that the map going from the currents in a 2d CFT to the thermal correlation functions of their zero-modes has a non-trivial kernel [^1] We would like to thank G. M. T. Watts for this observation.. That is, if we used the method previously mentioned, then we would not know a priori whether certain parts of that expression vanished.
@@
Given that the charges $I_{2n-1}(L)$ do not commute with the charges $J_{2n-1}(L)$, one question that may be asked is “Is this re-exponentiation a reasonable thing to do?". It seems that the answer is yes, and the fact that these charges do not commute does not affect our ability to formally re-write the transformed GGE as another GGE. Let us take some time to elaborate on this point. When we expand the right hand side of (57), the ordering of the charges in the correlators matters since they do not commute. This will lead to the presence of correlators that contain the same charges in different orders and we need to ensure that all the necessary correlators are present when we take the modular transformations of the $\langle{I_{5}^{n}}\rangle$ in the original GGE.
@@
It is worth mentioning briefly that $\langle{I_{5}J_{9}I_{5}J_{9}}\rangle$ and $\langle{I_{5}^{2}J_{9}^{2}}\rangle$ cannot independently be written as modular linear differential operators (MLDOs) acting on the characters of the theory, but this particular linear combination presented above does have a representation as an MLDO acting on the characters. We suspect that if one carefully studies the contact terms between the relevant currents associated to these charges, as was done in [13] for a different model, then it may become clear that indeed these expectation values separately cannot be written as MLDOs, however we have not performed this analysis.
@@
While we have found an asymptotic expression for the transformed GGE, or rather an expression with the leading charges in the transformed GGE, (57), we don’t believe that these match as functions. Firstly the right hand side of (57) contains an infinite sum in the charges. It is not clear if this sum is convergent, indeed in the case of free fermions the equivalent sum over charges was not convergent and had to be regularised [15]. In the case of free fermions this regularisation introduced functions with a branch cut. Hence while the original GGE was real, the transformed expression was complex. This problem was resolved by introducing additional terms in the transformed expression that came from the thermodynamic Bethe ansatz (TBA). These additional terms made the transformed expression real. It was then proved in [17] that these additional terms gave expressions that matched exactly, not just asymptotically. We will now use the TBA for the Lee-Yang model to first reproduce our asymptotic results, and then find additional terms that we believe should be included in the transformed expression for the GGE.